# Reinformer: Max-Return Sequence Modeling for Offline RL

### v2

@@ 2.2 Sequence Modeling in Reinforcement Learning
Compared to online reinforcement learning which interacts with the environment, offline RL is more similar to the data-driven paradigm given the offline dataset $\mathcal{D}$. As a result, offline RL (Chen et al., 2021) has been formulated as the supervised sequence modeling, different from the classical MDP formulation. The offline dataset can be denoted as the sequence form rather than transitions $\mathcal{D}=\left\{\left(\cdots,s^{\left(n\right)}_{t},a^{\left(n\right)}_{t},g^{\left(n\right)}_{t}\cdots\right)\right\}$. Here $g^{\left(n\right)}_{t}$ is the returns-to-go (or simply returns) defined as $g^{\left(n\right)}_{t}\dot{=}\sum_{t^{\prime}=t}^{T}r\left(s_{t^{\prime}}^{\left(n\right)},a_{t^{\prime}}^{\left(n\right)}\right)$ that represents the sum of future rewards from current timestep $t$.
Decision Transformer (DT) (Chen et al., 2021), following the upside-down RL (Srivastava et al., 2019; Schmidhuber, 2019), predicts the actions based on the previous trajectories $\tau$ concatenated with returns-to-go $g_{t}^{\left(n\right)}$:
@@ 3.1 Trajectory Stitching Example
In the offline RL literature, trajectory stitching receives lots of attention. Ideally, the offline agent should take suboptimal trajectories that overlap and stitch them into the optimal trajectories (Kostrikov et al., 2021b). It has been proved both theoretically (Brandfonbrener et al., 2022) and empirically (Kumar et al., 2022) that the return-conditioned sequence modeling lacks stitching ability. We utilize the following example to thoroughly describe this.
@@ 3.3 Implementation Details
with $\text{with \ }\Delta g=g_{t}^{(n)}-$ $\pi_{\theta}\left(\mathbf{S}^{\left(n\right)}_{t-K}\right)$
@@ 3.4 Comparison with EDT
Elastic Decision Transformer (EDT) (Wu et al., 2023) also explicitly considers trajectory stitching, so we briefly introduce EDT and then compare it with our Reinformer. EDT also introduces the same returns-to-go loss 4. During the inference phase, EDT uses this loss to estimate the maximum returns-to-go $\hat{g}_{\text{max}}(K)$ achievable under different historical lengths $K$. EDT then selects the maximum $\hat{g}^{*}_{\text{max}}=\max_{K}\hat{g}_{\text{max}}(K)$ and finally determine the action based on the Expert Action Inference (from Multi-game Decision Transformer (Lee et al., 2022)) along with the EDT model. The underlying idea behind EDT’s inference method is to dynamically adjust historical trajectories, preserving longer historical lengths when previous trajectories are optimal and maintaining shorter historical lengths when they are suboptimal.
@@
hopper-medium 52.9 58.5 66.3 67.6 67.0 63.5 66.5 80.02 $\pm$6.74
hopper-medium-replay 18.1 95.0 94.7 82.7 86.6 89.0 52.1 83.67 $\pm$5.26
@@ 4 Related Work
As the cradle of supervised sequence modeling, offline RL (Levine et al., 2020; Fujimoto et al., 2019) breaks free from the traditional paradigm of online interaction. Classic off-policy actor-critic algorithms (Degris et al., 2012; Konda & Tsitsiklis, 1999) encounter out-of-distribution (OOD) issues, where the value function tends to overestimate the out-of-distribution (OOD) state-action pairs (Levine et al., 2020; Fujimoto et al., 2019). Mainstream offline algorithms can be categorized into two classes. One is policy constraint, which constrains the learned policy stay close to the behavior policy based on different “distance” such as batch constrained (Fujimoto et al., 2019), KL divergence (Wu et al., 2019), MMD distance (Kumar et al., 2019) and MSE constraint (Fujimoto & Gu, 2021). The other is value regularization, which regularizes the value function to assign low values on OOD state-action pairs (Kumar et al., 2020; Kostrikov et al., 2021a; Bai et al., 2022). Some algorithms take a different approach, understanding and solving offline RL problems from the perspective of on-policy learning. R-BVE (Gulcehre et al., 2021) and Onestep RL (Brandfonbrener et al., 2021) both transform off-policy style offline algorithms (such as CRR (Wang et al., 2020), BCQ (Fujimoto et al., 2019), BRAC (Wu et al., 2019)) into on-policy style. Besides, BPPO (Zhuang et al., 2023) finds the online algorithm PPO (Schulman et al., 2017) can directly solve the offline problem duo the inherent conservatism of PPO. In contrast, Decision Transformer (DT) (Chen et al., 2021) directly maximizes the action likelihood, which opens up a new paradigm of supervised sequence modeling to address offline RL.
Supervised Learning in Reinforcement Learning
Before Decision Transformer (DT), upside-down reinforcement learning (Srivastava et al., 2019; Schmidhuber, 2019) has already begun exploring RL solutions using supervised learning methods. DT (Chen et al., 2021) incorporates return as part of the sequence to predict the optimal action. This paradigm breaks away from the classic RL paradigm, circumventing OOD problems directly. Inspired by DT, RL is investigated from a supervised learning perspective, including network architecture (Kim et al., 2023; David et al., 2022), unsupervised pretraining (Xie et al., 2023) and large capacity model (Lee et al., 2022). While other works improve DT using RL components such as online finetuning (Zheng et al., 2022), trajectory stitching (Wu et al., 2023) and dynamics programming (Yamagata et al., 2023). However, no sequence model incorporates the RL objective that maximizes returns to enhance the model.
@@ 5.1 Results on D4RL Benchmark
umaze 68.5 84.0 64.5 67.8 53.1 84.4±2.7
umaze-diverse 64.8 79.5 60.5 58.3 50.2 65.8±4.1
medium-play 4.5 78.5 0.8 0.0 0.0 13.2±6.1
medium-diverse 4.8 83.5 0.5 0.0 0.0 10.6±6.9
@@ Appendix B Experiment Details and Hyperparameters
The state tokens, return tokens and action tokens are first processed by different linear layers. Then these tokens are fed into the decoder layer to obtain the embedding with dimension 256. Here the decoder layer is a lightweight implementation from “min-decision-transformer”. The context length for the decoder layer is denoted as $K$. We use the LAMB (You et al., 2019) optimizer with to optimize the model with action loss and returns-to-go loss. The hyperparameter of returns loss is denoted as $m$. The parameters of different dataset are summarized as follow:
datasets $m$ train step layers heads K lr
halcheetah-medium 0.9 50000 4 8 5 0.0001
halcheetah-medium-replay 0.9 50000 4 8 5 0.0001
halcheetah-medium-expert 0.9 50000 4 8 20 0.0001
Hopper-medium 0.999 35000 4 8 5 0.0001
Hopper-medium-replay 0.999 200000 4 8 5 0.0001
Hopper-medium-expert 0.9 100000 4 8 20 0.0001
walker2d-medium 0.9 15000 4 8 5 0.0001
walker2d-medium-replay 0.99 80000 4 8 2 0.0004
walker2d-medium-expert 0.99 50000 4 8 20 0.0001
Maze2d-umaze 0.99 50000 3 8 20 0.0001
Maze2d-medium 0.99 35000 5 8 10 0.0004
Maze2d-large 0.99 45000 5 8 10 0.0004
Kitchen-complete 0.99 100000 5 8 20 0.0001
Kitchen-partial 0.9 90000 5 8 20 0.0001
Kitchen-mixed 0.9 50000 5 8 20 0.0001
Antmaze-umaze 0.9 50000 3 8 2 0.0001
Antmaze-umaze-diverse 0.99 50000 3 8 2 0.0001
Antmaze-medium-play 0.99 100000 6 8 3 0.0008
Antmaze-medium-diverse 0.99 100000 5 4 2 0.0004
Table 3: Hyperparameters.
Appendix C Training Curves on Gym-v2, Maze2d-v1 and Kitchen-v0
We exhibit the training curves on five seeds with the mean and the standard deviation.
                       (a) Maze2d-umaze                               (b) Maze2d-medium                                 (c) Maze2d-large
Figure 8: The training curves on Maze2d-v1.
                      (a) Kitchen-complete                            (b) Kitchen-partial                          (c) Kitchen-mixed
Figure 9: The training curves on Kitchen-v0.
                  (a) HalfCheeth-medium                    (b) HalfCheeth-medium-replay               (c) HalfCheeth-medium-expert

### v3

@@ 2.2 Sequence Modeling in Reinforcement Learning
Compared to online reinforcement learning which interacts with the environment, offline RL is more similar to the data-driven paradigm given the offline dataset $\mathcal{D}$. As a result, offline RL (Chen et al., 2021) has been formulated as the supervised sequence modeling, different from the classical MDP formulation. The offline dataset can be denoted as the sequence form rather than transitions $\mathcal{D}=\left\{\left(\cdots,s^{\left(n\right)}_{t},a^{\left(n\right)}_{t},g^{\left(n\right)}_{t}\cdots\right)\right\}$. Here $g^{\left(n\right)}_{t}$ is the returns-to-go (or simply returns) defined as $g^{\left(n\right)}_{t}\dot{=}\sum_{t^{\prime}=t}^{T}r\left(s_{t^{\prime}}^{\left(n\right)},a_{t^{\prime}}^{\left(n\right)}\right)$ that represents the sum of future rewards from current timestep $t$. Decision Transformer (DT) (Chen et al., 2021), following the upside-down RL (Srivastava et al., 2019; Schmidhuber, 2019), predicts the actions based on the previous trajectories $\tau$ concatenated with returns-to-go $g_{t}^{\left(n\right)}$:
@@ 3.1 Trajectory Stitching Example
In the offline RL literature, trajectory stitching receives lots of attention. Ideally, the offline agent should take suboptimal trajectories that overlap and stitch them into the optimal trajectories (Kostrikov et al., 2021b; Liu et al., 2023). It has been proved both theoretically (Brandfonbrener et al., 2022) and empirically (Kumar et al., 2022) that the return-conditioned sequence modeling lacks stitching ability. We utilize the following example to thoroughly describe this.
@@ 3.3 Implementation Details
with $\text{with \ }\Delta g=g_{t}^{(n)}-$ $\pi_{\theta}\left(\mathbf{S}^{\left(n\right)}_{t-K}\right).$
@@ 3.4 Comparison with EDT
Elastic Decision Transformer (EDT) (Wu et al., 2023) explicitly considers trajectory stitching, so we briefly introduce EDT and then compare it with our Reinformer. EDT also introduces the same returns-to-go loss (4). During the inference phase, EDT uses this loss to estimate the maximum returns-to-go $\hat{g}_{\text{max}}(K)$ achievable under different historical lengths $K$. EDT then selects the maximum $\hat{g}^{*}_{\text{max}}=\max_{K}\hat{g}_{\text{max}}(K)$ and finally determine the action based on the expert action inference (Lee et al., 2022) along with the EDT model. The inference of EDT dynamically adjusts historical trajectories, which preserves longer historical lengths when previous trajectories are optimal and shortens them when they are suboptimal.
@@
hopper-medium 52.9 58.5 66.3 67.6 67.0 63.5 66.5 81.60 $\pm$3.32
hopper-medium-replay 18.1 95.0 94.7 82.7 86.6 89.0 52.1 83.31 $\pm$3.47
@@ 4 Related Work
As the cradle of supervised sequence modeling, offline RL (Levine et al., 2020) breaks free from the traditional paradigm of online interaction. Classic off-policy actor-critic algorithms (Degris et al., 2012; Konda & Tsitsiklis, 1999) encounter out-of-distribution (OOD) issues, where the value function tends to overestimate the OOD state-action pairs (Fujimoto et al., 2019; Liu et al., 2024c). Mainstream offline algorithms can be categorized into two classes. One is policy constraint, which constrains the learned policy stay close to the behavior policy based on different “distance” such as batch constrained (Fujimoto et al., 2019), KL divergence (Wu et al., 2019; Liu et al., 2022), MMD distance (Kumar et al., 2019) and MSE constraint (Fujimoto & Gu, 2021). The other is value regularization, which regularizes the value function to assign low values on OOD state-action pairs (Liu et al., 2024b; Kostrikov et al., 2021a; Bai et al., 2022). Some algorithms take a different approach, understanding and solving offline RL problems from the perspective of on-policy learning. R-BVE (Gulcehre et al., 2021) and Onestep RL (Brandfonbrener et al., 2021) both transform off-policy style offline algorithms (such as CRR (Wang et al., 2020), BCQ (Fujimoto et al., 2019), BRAC (Wu et al., 2019)) into on-policy style. Besides, BPPO (Zhuang et al., 2023) finds the online algorithm PPO (Schulman et al., 2017) can directly solve the offline RL due to its inherent conservatism. In contrast, Decision Transformer (DT) (Chen et al., 2021) directly maximizes the action likelihood, which opens up a new paradigm called supervised sequence modeling.
Sequence Modeling in Reinforcement Learning
Before Decision Transformer (DT), upside-down reinforcement learning (Srivastava et al., 2019; Schmidhuber, 2019) has already begun exploring RL solutions using supervised learning methods. DT (Chen et al., 2021) incorporates return as part of the sequence to predict the optimal action. This paradigm breaks away from the classic RL paradigm, circumventing OOD problems directly. Inspired by DT, RL is investigated from a supervised learning perspective, including network architecture (Kim et al., 2023; David et al., 2022), unsupervised pretraining (Xie et al., 2023) and large capacity model (Lee et al., 2022). While other works improve DT using RL components such as online finetuning (Zheng et al., 2022), trajectory stitching (Wu et al., 2023) and dynamics programming (Yamagata et al., 2023). However, no sequence model incorporates the RL objective that maximizes returns to enhance the model (Liu et al., 2024a).
@@ 5.1 Results on D4RL Benchmark
umaze 68.5 84.0 64.5 67.8 53.1 84.4$\pm$2.7
umaze-diverse 64.8 79.5 60.5 58.3 50.2 65.8$\pm$4.1
medium-play 4.5 78.5 0.8 0.0 0.0 13.2$\pm$6.1
medium-diverse 4.8 83.5 0.5 0.0 0.0 10.6$\pm$6.9
@@ Appendix B Experiment Details and Hyperparameters
The state tokens, return tokens and action tokens are first processed by different linear layers. Then these tokens are fed into the decoder layer to obtain the embedding. Here the decoder layer is a lightweight implementation from “min-decision-transformer”. The context length for the decoder layer is denoted as $K$. We use the LAMB (You et al., 2019) optimizer with to optimize the model with action loss and returns-to-go loss. The hyperparameter of returns loss is denoted as $m$.
B.1 Hyperparameter $m$
The hyperparameter $m$ is crucially related to the returns-to-go loss and is one of our primary focuses for tuning. We explore values within the range of $m=[0.7,0.9,0.99,0.999]$. When $m=0.5$, the expectile loss function will degenerate into MSE loss, which means the model is unable to output a maximized returns-to-go. So we do not take $m=0.5$ into consideration. We observe that performance is generally lower at $m=0.7$ compared to others. Only hopper-medium and hopper-medium-replay adopt the parameter $m=0.999$ while $m=0.9$ and $m=0.99$ are generally better than $m=0.999$ on other datasets. The detailed hyperparameter selection of $m$ is summarized in the following table:
Table 3: Hyperparameters $m$ of returns loss on different datasets.
Dataset $m$ maze2d-umaze 0.99
halfcheetah-medium 0.9 maze2d-medium 0.99
halfcheetah-medium-replay 0.9 maze2d-large 0.99
halfcheetah-medium-expert 0.9 kitchen-complete 0.99
hopper-medium 0.999 kitchen-partial 0.9
hopper-medium-replay 0.999 kitchen-mixed 0.9
hopper-medium-expert 0.9 Antmaze-umaze 0.9
walker2d-medium 0.9 Antmaze-umaze-diverse 0.99
walker2d-medium-replay 0.99 Antmaze-medium-play 0.99
walker2d-medium-expert 0.99 Antmaze-medium-diverse 0.99
B.2 Context Length $K$
The context length $K$ is another key hyperparameter for sequence modeling, and we conduct a parameter search across the values $K=[2,5,10,20]$. The maximum value is $20$ because the default context length for DT (Chen et al., 2021) is $20$. The minimum is $2$, which corresponds to the shortest sequence length (setting $K=1$ would no longer constitute sequence modeling). Overall, we found that $K=20$ leads to more stable learning and better performance on high quality dataset (such as Gym-medium-expert and Kitchen). Conversely, a smaller context length is preferable on low quality dataset (such as Gym-medium/medium-replay and Antmaze). The parameter $K$ has been summarized as follows:
Table 4: Context length $K$ on different datasets.
Dataset $K$ maze2d-umaze 20
halfcheetah-medium 5 maze2d-medium 10
halfcheetah-medium-replay 5 maze2d-large 10
halfcheetah-medium-expert 20 kitchen-complete 20
hopper-medium 5 kitchen-partial 20
hopper-medium-replay 5 kitchen-mixed 20
hopper-medium-expert 20 Antmaze-umaze 2
walker2d-medium 5 Antmaze-umaze-diverse 2
walker2d-medium-replay 2 Antmaze-medium-play 3
walker2d-medium-expert 20 Antmaze-medium-diverse 2
B.3 Training Steps and Learning Rate
The default number of training steps is 50000, with a learning rate of 0.0001. With these default settings, if the training score continues to rise, we would consider increasing the number of training steps or doubling the learning rate. For some datasets, 50000 steps may cause overfitting and less training steps are better. The training steps are presented in Table 5 and learning rate is summarized in Table 6. We evaluate the policy every 5000 steps to obtain a normalized score. For each seed, this normalized score is calculated as the average returns of 100 trajectories except for 10 trajectories on Gym-v2 datasets.
Table 5: The training steps on different datasets.
Dataset Training Steps maze2d-umaze 50000
halfcheetah-medium 50000 maze2d-medium 35000
halfcheetah-medium-replay 50000 maze2d-large 45000
halfcheetah-medium-expert 50000 kitchen-complete 100000
hopper-medium 30000 kitchen-partial 90000
hopper-medium-replay 300000 kitchen-mixed 50000
hopper-medium-expert 100000 Antmaze-umaze 50000
walker2d-medium 15000 Antmaze-umaze-diverse 50000
walker2d-medium-replay 80000 Antmaze-medium-play 100000
walker2d-medium-expert 50000 Antmaze-medium-diverse 100000
Table 6: The learning rate on different datasets.
Learning Rate Dataset
0.0008 Antmaze-medium-play;
0.0004 1) walker2d-medium-replay;
2) maze2d-medium; 3) maze2d-large;
4) Antmaze-medium-diverse;
0.0001 Other datasets.
B.4 Network Architecture
The default values for the number of decoder layer, attention heads and hidden dimension are 4, 8 and 256, respectively. These parameters are usually fixed. When we observe an initial increase followed by a decrease in the training curve, we infer overfitting and reduce the number of layers. On the contrary, if the training curve consistently rises without a clear convergence trend, we would attempt to increase the number of layers. As for the number of attention heads, only Antmaze-medium-diverse is 4.
Table 7: The number of encoder layers and attention heads on different datasets.
Hyperparameters Values Dataset
Encoder Layers 3 maze2d-umaze; antmaze-umaze; antmaze-umaze-diverse;
4 Gym-v2;
5 maze2d-medium; maze2d-large; Kitchen-v0; Antmaze-medium-diverse;
6 Antmaze-medium-play.
Attention Heads 4 Antmaze-medium-diverse;
8 Other datasets.
Appendix C Training Curves
We exhibit the training curves on five seeds. The black line represents the mean of these five seeds and the red shaded area represents the variance.
C.1 Gym-v2
The training curves on nine datasets from Gym-v2 are plotted in Figure 8. Among these nine datasets, hopper-medium-replay exhibits severe training fluctuations, while walker2d-medium-replay shows slight fluctuation. The remaining datasets are notably stable, yielding satisfactory results without deliberate hyperparameter tuning.
             (a) halfcheetah-medium     (b) halfcheetah-medium-replay    (c) halfcheetah-medium-expert
@@ C.1 Gym-v2
C.2 Maze2d-v1
The training curves on three datasets from Maze2d-v1 are plotted in Figure 9. The training curve of maze2d-umaze is relatively stable. The variance on dataset maze2d-medium is very high and the training process also suffers from severe training fluctuations. Sometimes, the score can even approach to 125. Dataset maze2d-medium also fluctuates a little.
                    (a) maze2d-umaze                            (b) maze2d-medium                             (c) maze2d-large
Figure 9: The training curves on Maze2d-v1.
C.3 Kitchen-v0
The training curves on three datasets from Kitchen-v0 are plotted in Figure 9. Overall, the training curves are remarkably stable, and the long context length $K$ is crucial for the stability of learning.
               (a) kitchen-complete                     (b) kitchen-partial                       (c) kitchen-mixed
Figure 10: The training curves on Kitchen-v0.
C.4 Antmaze-v2
Since we report the best score during training rather than the last score, we do not provide training curves on Antmaze. Here, we would like to emphasize the reward modification. Antmaze contains datasets with sparse rewards, where only 1 indicates the reach of the goal while 0 is not. In order to enhance the reward signal, we multiply the reward by 100. However, we found that this modification leads to the occurrence of NaN values on dataset Antmaze-umaze-diverse. Besides, the original reward also occur the NaN values. So we modify the reward by adding another 1, that is, $\hat{r}=100\times r+1$. We summarize the results of different reward modification in Table 8.
Table 8: The normalized best score on Antmaze-v2 dataset with different reward modification.
Antmaze $\hat{r}=100\times r$ $\hat{r}=100\times r+1$
umaze 85$\pm$3.87 84.4$\pm$2.7
umaze-diverse NaN 65.8$\pm$4.1
medium-play 11.4$\pm$3.78 13.2$\pm$6.1
medium-diverse 7.2$\pm$2.17 10.6$\pm$6.9